150619 SFSU

We got back up and running Weds 0617 with a new power supply – evening looked promising and fan was on but no sample next morning.

Thurs 0618 installed previous SD card triggering when RH>89% trigger. Drier evening, higher cloud, no trigger.

Today, Fri 0619 we installed the passive sensor, which is currently running (as of 1400 PDT Daniel), though not wired to active sampler yet. Another breakthrough. Have found access to ethernet, will install Tuesday next week.

Forecast for warm, dry weekend, so the sampler is decommissioned until Monday 0622.

Leave a Reply

Your email address will not be published. Required fields are marked *